OpenStudy (anonymous):

Find the slope intercept form for the line satisfying the following conditions. x intercept 3, y intercept 1/3 what is the slope intercept form for the line?

OpenStudy (anonymous):

(3,0) and (0,1/3) so the slope is (-1/3)/3=-1/9 y=-(1/9)x+1/3